OpenStudy (anonymous):

What Is 6 Rounded To The Nearest Tenths?

OpenStudy (anonymous):

0

OpenStudy (campbell_st):

it would be 60 tenths or 1/10 are the 1st decimal place. and 1 = 10/10 so 6 = 60/10